Rad Co. provides the following sales forecast and production budget for the next four months: April May June July 500 580 530 600 Budgeted production (units) 442 570 544 540 Sales (units) The company plans for finished goods inventory of 120 units at the end of June. In addition, each finished unit requires 5 pounds of raw materials at cost of $2.00 per pound and the company wants to end each month with raw materials inventory equal to 30% of next months production needs. Beginning raw materials inventory for April was 663 pounds. Each finished unit requires 0.50 hours of direct labor at the rate of $16 per hour. The company budgets variable overhead at the rate of $20 per direct labor hour and budgets fixed overhead of $8,000 per month. Prepare a raw materials budget for April, May, and June. RAD CO. Direct Materials Budget For April, May, and June April May June Budget production (units) Budgeted ending inventory (Ibs.) Materials needed for production (Ibs.) 442 570 544 units 570 544 540 lbs. Total materials requirements (lbs.) Materials to be purchased (Ibs.) Materials price per pound Total cost of direct materials purchases 0 0
Answer #1

Solution :-

RAD CO .

Direct materials budget

For April, May and June

April May June
Budget production ( unit ) 442 570 544
Materials requirement per unit 5 5 5
Materials needed for production ( lbs )

= 442 * 5

= 2,210

= 570 * 5

= 2,850

= 544 * 5

= 2720

Budgeted ending inventory at 30% of next month

= 2,850 * 30%

= 855

= 2,720 * 30%

= 816

= [ 540 * 5 ] * 30%

= 810

Total material requirement

= 2,210 +  855

= 3,065

= 2,850 + 816

= 3,666

= 2720 + 810

= 3,530

Beginning inventory ( lbs ) 663 855 816
Materials to be purchased

= 3,065 - 663

= 2,402

= 3,666 - 855

= 2,811

= 3,530 - 816

= 2,714

Material price per pound $20 $20 $20
Budgeted cost of direct materials purchases

= 2,402 * $20

= $48,040

= 2,811 * $20

= $56,220

= 2,714 * $20

= $54,280